OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (texaschic101):

first we find the slope by using our points and the slope formula slope(m) = (y2 - y1) / (x2 - x1) (0,0)...x1 = 0 and y1 = 0 (4,-4)...x2 = 4 and y2 = -4 now we sub slope(m) = (-4-0) / (4 - 0) slope(m) = -4/4 slope(m) = -1 now we use y = mx + b slope(m) = -1 use either of your sets of points..(4,-4)...x = 4 and y = -4 now we sub -4 = -1(4) + b -4 = -4 + b -4 + 4 = b 0 = b so your equation is : y = -x + 0 standard form is Ax + By = C add x to both sides x + y = 0 <=== standard form

OpenStudy (texaschic101):

any questions ?